OpenStudy (anonymous):

Factor by factoring out the greatest common factor x^2-5x

OpenStudy (anonymous):

\[x = 0\] \[x - 5 = 0 \rightarrow x - 5 + 5 = 0 + 5\] \[x = 5\]

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

no problem :)